Swifty Company purchases equipment on January 1, Year 1, at a cost of $486,000. The asset is expected to have a service life of 12 years and a salvage value of $43,740.

Collapse question part

(a)

Correct answer. Your answer is correct.
Compute the amount of depreciation for each of Years 1 through 3 using the straight-line depreciation method. (Round answers to 0 decimal places, e.g. 5,125.)

Depreciation for Year 1

$enter a dollar amount rounded to 0 decimal places

Entry field with correct answer

Depreciation for Year 2

$enter a dollar amount rounded to 0 decimal places

Entry field with correct answer

Depreciation for Year 3

$enter a dollar amount rounded to 0 decimal places

Entry field with correct answer

SHOW SOLUTION

SHOW ANSWER

LINK TO TEXT

Attempts: 1 of 3 used

Collapse question part

(b)

Correct answer. Your answer is correct.
Compute the amount of depreciation for each of Years 1 through 3 using the sum-of-the-years'-digits method.

Depreciation for Year 1

$enter a dollar amount

Entry field with correct answer

Depreciation for Year 2

$enter a dollar amount

Entry field with correct answer

Depreciation for Year 3

$enter a dollar amount

Entry field with correct answer

SHOW SOLUTION

SHOW ANSWER

LINK TO TEXT

Attempts: 1 of 3 used

Collapse question part

(c)

Compute the amount of depreciation for each of Years 1 through 3 using the double-declining-balance method. (Round depreciation rate to 2 decimal places, e.g. 15.84%. Round answers to 0 decimal places, e.g. 45,892.)

Depreciation for Year 1

$enter a dollar amount rounded to 0 decimal places

Depreciation for Year 2

$enter a dollar amount rounded to 0 decimal places

Depreciation for Year 3

$enter a dollar amount rounded to 0 decimal places
Answer #1
a
Depreciation for Year 1 36855 =(486000-43740)/12
Depreciation for Year 2 36855 =(486000-43740)/12
Depreciation for Year 3 36855 =(486000-43740)/12
b
Sum-of-the-years'-digits 78 =12*(12+1)/2
Depreciation for Year 1 68040 =(486000-43740)/78*12
Depreciation for Year 2 62370 =(486000-43740)/78*11
Depreciation for Year 3 56700 =(486000-43740)/78*10
c
Depreciation rate 16.67% =1/12*2
Depreciation for Year 1 81016 =486000*16.67%
Depreciation for Year 2 67511 =(486000-81016)*16.67%
Depreciation for Year 3 56257 =(486000-81016-67511)*16.67%